OpenStudy (anonymous):

will award metal if right answer Which line is not perpendicular to the line 4x − 3y = 2? 3x + 4y = 2 3x − 4y = 2 -3x − 4y = 5 3x + 4y = -8

OpenStudy (mathstudent55):

What do you know about the slopes of perpendicular lines?

OpenStudy (anonymous):

nothing thats why i put it up there

OpenStudy (anonymous):

3 + 3 =9 sooo i think is B

OpenStudy (anonymous):

perpendicular lines make a right angle when they intersect

OpenStudy (anonymous):

so wats da answer n nice name

OpenStudy (mathstudent55):

The slopes of perpendicular lines are negative reciprocals. That means that when you multiply together the slopes of two perpendicular lines, you get -1.

OpenStudy (mathstudent55):

Here is the equation of a line in standard form: \(ax + by = k\) We solve it for y to get it in slope-intercept form. \(by = -ax + k\) \(y = -\dfrac{a}{b}x + \dfrac{k}{b} \) The slope-intercept form is \(y = mx + b\), where m is the slope. That means, to find the slope using the standard form, just find \(-\dfrac{a}{b} \). Take the given equation and each of the choices, and find \(-\dfrac{a}{b} \). Then find which choice has a slope that is not the negative reciprocal of the slope of the given equation.